OpenStudy (anonymous):

find the polynominal equation of least degree with roots -3,0, and 3

OpenStudy (anonymous):

i got x^3+9x=0

Directrix (directrix):

The polynomial would be written as P(x) = x^3 + 9x. I don't know if it is correct but I will see.

Directrix (directrix):

P(x) = x^3 + 9x. P(X) = 0 0 = x^3 + 9x 0 = x(x^2 + 9) 0 = x or x^2 + 9 = 0 x = 0 or x^2 = -9 x = 0 or x = 3i or -3i

Directrix (directrix):

NOT Correct: i got x^3+9x=0 Try again. @Justmeneedinghelp

Directrix (directrix):

So, what is the correct answer?